OpenStudy (anonymous):

-x + y =3 and 2x-2y=-6

OpenStudy (anonymous):

what are you trying to solve? make it in y= form or what?

OpenStudy (anonymous):

its sytems. it says to solve and explain what the solution tells us about the lines represented by equations

OpenStudy (anonymous):

they are the same line... watch if i mutliply the first equation by -2 -2(-x+y)=-2*3 2x-2y=-6 which is the same as the second line

OpenStudy (anonymous):

ax+by=c is called standard form of the line, it has some posotive atributes but it is not unique; the slope-intercept form of a line(y=mx+b) is unique and you would see that they are the same line if you solved each for y.

OpenStudy (anonymous):

Depending on what your teacher/text wants is how you indicate the solution, but geometrically the lines are COINCIDENT (on top of each other).

OpenStudy (anonymous):

The solution set often looks like {(x,y):y=x+3}

OpenStudy (anonymous):

thanks my teacher did one with us and he didnt explain

OpenStudy (anonymous):

ru good

OpenStudy (anonymous):

yea thanks